OpenStudy (anonymous):

how do I evaluate 2^3 / 2^0 - 2^-1?

OpenStudy (anonymous):

7 1/2? cause 8 divided by 1 is 8 subtracted from 1/2 which is 2^-1?

OpenStudy (anonymous):

(8/1) -(1/2) = 15/2

OpenStudy (anonymous):

steps please :(

OpenStudy (anonymous):

2 to the 3rd power is 8, and 2^0 is 1 and 2^-1 is 1/2

OpenStudy (anonymous):

\[(2{}^{\wedge}3/2{}^{\wedge}0)-(2{}^{\wedge}-1) \]\[(8)-\left(\frac{1}{2}\right) \]\[\frac{15}{2} \]

OpenStudy (anonymous):